OpenStudy (anonymous):

If r = 45 cm and q = 53 cm, what is the length of p?

OpenStudy (anonymous):

A 28 cm B 26 cm C 70 cm

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

\[r^{2} + p^{2} = q^{2}\]

OpenStudy (ipwnbunnies):

You should know the pytharogean theorem by now based on the amount of time you're around here lol.

OpenStudy (anonymous):

lol nope

OpenStudy (ipwnbunnies):

You don't know the Pythagorean Theorem at all?

OpenStudy (anonymous):

I do but its confusing I don't know how to do it

OpenStudy (ipwnbunnies):

Ok, bear with me, it only applies to right triangles. Right triangles are the triangles with 90 degree angles.

OpenStudy (ipwnbunnies):



OpenStudy (anonymous):

ok

OpenStudy (ipwnbunnies):

So, what the theorem is getting at is that the square of the shortest side + the square of the longer side = the square of the longest side

OpenStudy (ipwnbunnies):

a*a + b*b = c*c

OpenStudy (anonymous):

yeah I get that part

OpenStudy (ipwnbunnies):

Good, so let's put it in the terms of your problem: p*p + r*r = q*q

OpenStudy (anonymous):

yeah but don't we fill in the numbers

OpenStudy (ipwnbunnies):

We are given the length of the longest side, q, and the length of the middle side, r. Solve for p.

OpenStudy (ipwnbunnies):

Yeah, we're getting to the numbers. Just follow my instructions and solve for p in the pythagorean theorem.